# Phase diagram of the quantum spin-1/2 Heisenberg-$\Gamma$ model on a frustrated zigzag chain

### v1

@@ II.2.2 Correlation functions
We obtain the two-point correlation functions $\langle O_{i}O_{j}\rangle$, using the single-site or two-site operator $O_{i}$. For $O_{i}$, we take spin operators, $S_{i}^{\alpha}$ with $\alpha=x,y,z$, vector chiral operator $\kappa_{i}^{(n)}=S_{i}^{x}S_{i+n}^{y}-S_{i}^{y}S_{i+n}^{x}$, and nematic operator $Q^{xy}_{i}=S_{i}^{x}S_{i+1}^{y}+S_{i}^{y}S_{i+1}$. Figure 4 shows these correlation functions as functions of distance, $|i-j|$, obtained using SSD-DMRG. Let us summarize the features of each phase.
@@
TLL phase. Basically the correlation functions all decay algebraically up to $|i-j|\leq 50$, where we reach half of the system size. However, at $\theta\neq 0$, $\langle Q_{i}^{xy}Q_{j}^{xy}\rangle$ shows a robust constant value at large distances, unlike the standard TLL phase. This is because if the system remains paramagnetic without any other orderings, the $\Gamma$ term works as a conjugate field to induce a finite value of $\langle Q_{i}^{xy}\rangle$, which is obvious from Eq.(1). In that context, the emergent $Q_{i}^{xy}$ is trivial as it does not break any symmetry of the Hamiltonian.
@@
Nematic-singlet dimer (ND) phase. Along the exactly solvable MG line $J_{2}/J_{1}=\Gamma_{2}/\Gamma_{1}=1/2$ at $\theta<{\rm arctan}(\sqrt{3})$, we find the dimer-product singlet states as an exact ground state, which strictly excludes other components and gives $\langle Q_{i}^{xy}Q_{j}^{xy}\rangle\rightarrow 0$. While, even when we are away from the MG line, the ND phase sustains where the magnetic correlation functions all decay exponentially, indicating te existence of a spin gap. Such gap opening is due to the breaking of translational symmetry in the same manner as the MG solution. The nematic correlation $\langle Q_{i}^{xy}Q_{j}^{xy}\rangle$ starts to converge to a small but finite constant value at long enough distances, indicating the formation of a long-range order with $\Gamma_{\eta}\neq 0$, as shown in Fig. 4, whose implication will be discussed shortly.
@@
Vector chiral (VC) phase. In the F-AF Heisenberg interaction range of the phase diagram at $\phi/\pi\sim 0.7-0.8$, we find a VC phase where the magnetic order is absent and $\langle\kappa^{(1)}_{i}\kappa^{(1)}_{j}\rangle$ sustains at long distances. Its amplitude shows a large oscillation in the period of chirality over about $30$ spins as can be seen from the behavior of $\langle S_{i}^{x}S_{j}^{x}\rangle$ and $\langle S_{i}^{x}S_{j}^{y}\rangle$, while the pure magnetic components decay in power law.
@@
Magnetically ordered phases. The F phases, AF-UD, and UUDD phases all have their magnetic easy axis pointing in the $x^{\prime}$ or $y^{\prime}$ directions perpendicular to the leg. Indeed, the correlation functions show a clear exponential decay about $\langle S^{z}_{i}S^{z}_{j}\rangle$, while those of the in-plane elements robustly take the constant value of order-1 throughout $|i-j|$. The nematic correlation is also robust because these magnetic orderings are the condensation of the off-diagonal $S=1$ elements.
II.2.3 Spin gap
@@
Our model does not conserve total-$S^{z}$, in which case the spin gap cannot be evaluated by the standard treatment of measuring the lowest-energy difference between different total-$S^{z}$ sectors. Instead, we apply a grand canonical appoach using SSDHotta and Shibata (2012); Hotta et al. (2013) that allows the access to the bulk magnetization curve for $N\gtrsim 20$ calculations (while we adopt $N=100$). By adding a Zeeman field $-h\sum_{j}S^{z}_{j}$ to Eq.(1), and by deforming them with the sine-square function, the magnetization density $M$ is obtained by extracting the intrinsic values near the center using the $q=0$ element of the SSD Fourier transformationKawano and Hotta (2022). Figure 5(a) shows four different magnetization curves obtained for the nonmagnetic ground states, TLL, ND, VC, and N. Only the ND phase shows a substantially large spin gap, $\Delta\sim 0.3$. The spin gap of $J_{1}-J_{2}$ chain ($\Gamma=0$) was previously calculated by DMRGWhite and Affleck (1996) and was evaluated as $\Delta=0.11\sim 0.37$ for $0.14\pi<\phi<0.25\pi$. For TLL and VC, the standard magnons condense and form a standard magnetization curve. Whereas, the N phase shows a rather unusual gapless ferro-like behavior of magnetization, which shall be because the nematic $S=1$ and $S^{z}=0$ component is precondensed in the ground state and the magnetic components have a close in energy.
@@
To confirm the presence of a finite spin gap, we examined the field-dependent magnetization in the direction of $-h\sum_{j}S^{\alpha}_{j}$, where $\alpha=x,y$, and $x^{\prime},y^{\prime}$ ($\pi/4$-rotation of $x,y$ about the $z$-axis) as shown in Appendix C; in the ND phase the spin gap is finite and does not depend much on the direction of a field.
II.2.4 Nematic order parameter
@@
We have shown in Fig. 4 that the nematic correlation, $\langle Q^{xy}_{i}Q^{xy}_{j}\rangle$, starts to converge to a robust constant value once we introduce finite $\Gamma/J\neq 0$. To examine how they behave at around $\Gamma/J\sim 0$, we plot in Fig. 5(b) the averages of $\overline{\langle Q^{xy}_{i}Q^{xy}_{j}\rangle}$ over $|i-j|=58-61$ to exclude the strong oscillation effect. Numerically, there is no “gap” in the onset value of $\big{(}\overline{\langle Q^{xy}_{i}Q^{xy}_{j}\rangle}\big{)}^{1/2}$, i.e. it increases immediately from $\Gamma=0$ in power of $\Gamma$ for all displayed parameters of $\phi$. As mentioned, $\Gamma$ in Eq.(1) is conjugate to $Q^{xy}_{i}=S_{i}^{x}S_{i+1}^{y}+S_{i}^{y}S_{i+1}^{x}$, that works as chemical potential or a Zeeman field of this nonmagnetic quasi-particle.
@@
The AF-UUDD has the staggard magnetic moment pointing in the $y^{\prime}$ direction which is nothing but the “magnetization” represented by $Q^{xy}$. Namely, the doped quasi-particles condense and form a regular four-fold periodic structure breaking the translational symmetry. In the ND phase, the two-fold periodic breaking of translational symmetry occurs but there is no magnetic order because of the interplay of singlet and nematic particles. The spin gap is open, which is distinct from the AF-UUDD and TLL phases.
@@
For these reasons, we understand that “field” effect, $\Gamma$, generates a finite nematic correlation but it appears in different ways depending on the degree of frustration dictated by $\phi$. Based on the bond-operator approach in the next section, we conclude that the nematic particles do not really condense in the ND phase but remain as fluctuations, while they finally condense in the magnetically ordered phases. The TLL phase, although hosting finite nematic correlation, shall be trivial.
@@ III Bond-operator mean-field theory
In this section, we examine the effect of $\Gamma$-term in the AF-AF zigzag chain ($J_{1},J_{2}>0,0\leq\phi\leq\pi/2$) by using bond-operator mean-field theorySachdev and Bhatt (1990). There are five independent nematic (quadrupolar) operators defined on a spin-1/2 pairs forming spin-1 given as Andreev and Grishchuk (1984)
In Eq.(1), the parameter $\Gamma$ is coupled to the $xy$ component $Q_{i}^{xy}=S_{i}^{x}S_{i+1}^{y}+S_{i}^{y}S_{i+1}^{x}$, and it works as a “field” (or a chemical potential) to condense the nematic particle represented by $Q_{i}^{xy}$. This situation is similar to doping a magnon to the gapped singlet stateSachdev and Bhatt (1990). Similarly, doping a nematic particle other than $xy$-component is done by a ring exchange interaction that works as a chemical potentialUeda and Totsuka (2007); Totsuka et al. (2012); Yokoyama and Hotta (2018). All these cases have a gapped singlet ground state as a starting point, approximated by the product state of the dimer (or tetramer) units, and the effect of the applied “field” is examined using the bond-operator approach or its analog.
@@ III.2 Effect of inter-dimer interaction
$S_{2}^{\alpha}$ $=\frac{s1}{2}(-s^{\dagger}p_{\alpha,n}-p_{\alpha,n}^{\dagger}s-i\epsilon_{\alpha\beta\gamma}p_{\beta,n}^{\dagger}p_{\gamma,n}),$ (7)
@@
Here, we consider a small $\lambda$, assuming $\gamma<2$, expecting that the ground state is dominated by singlets and the population of $p_{\alpha,n}$-bosons are dilute. The condensed singlets are replaced by its expectation value $\langle s\rangle=\bar{s}$, and by dropping off the terms consisting of three or four $p$ operators, we obtain the interactions between the $n$-th and $(n+1)$-th dimers as
@@ III.4 Phase diagram
The phase diagrams obtained by the bond-operator mean-field theory are shown in Figs. 6(b) and 6(c) for several fixed values of $\lambda=0.1,0.5,1.0$ and $\phi/\pi=0.08$ to $0.4$, respectively. The corresponding set of mean-field parameters $(\mu,\bar{s}^{2})$ are shown in Fig. 7 at $\lambda=1$ as functions of $\theta$. For these parameters, there is a range of small $\theta$ that the ground state is fully occupied by singlets, $\bar{s}^{2}=1$, which we call S-phase. The phase transition from S-phase to (S+N)-phase takes place at finite $\theta$, where $\bar{s}^{2}$ starts to decrease from 1, and the chemical potential $\mu$ reaches the singlet (zero) energy level. When increasing $\lambda$, the S-phase shrinks because the $p$-particles gain the kinetic energy due to inter-dimer interactions, but at $\phi={\rm atan}(1/2)$, i.e. $J_{2}/J_{1}=\Gamma_{2}/\Gamma_{1}=0.5$, which corresponds to the Majumdar-Ghosh line, the singlet phase remains stable throughout the whole range of $\theta$. This fact is trivial as we find that the inter-dimer interaction in Eq.(8) is $\hat{h}_{n}^{\rm MF}=0$, or equivalently, the dispersion in Eq.(12) yields $\epsilon(k)=0$ for $J_{1}=2J_{2}$. Indeed, along this line, the singlet product state is the exact solution of the original Hamiltonian Eq.(1) without any approximation.
@@
We plot in Figs. 8(a)-8(d) the energy dispersions $\omega_{l}(k)$ at $\lambda=1$ for several choices of $\phi$ and $\theta$, inside the S-phase. Remarkably, all of them are gapless. Indeed, the energy gap $\Delta$ of ${\cal H}_{\rm bo}$ shown in Fig. 8(e) is zero for all parameters of $\phi$ at small $\theta$. This shows that the singlet phase is robust but at the same time, the $p$-particle are always able to join the ground state. This result is rather unusual; the standard bond operator approach is used to provide the instability toward the condensation of these excited $p$-state that replaces the ground state, and the gap-closing point usually indicates the phase transition. Whereas, in the present case, the zero gap of ${\cal H}_{\rm bo}$ does not mean the instability of the singlet ground state. We emphasize that the instability takes place not at the gap closing point but when $\bar{s}^{2}$ starts to deviate from 1, and the chemical potential $\mu$ reaches the singlet energy level. For clarification, we plot in Fig. 7(b) together the bare excited energy, $\epsilon_{p_{-1}}$, of the isolated dimer. We see that $\mu$ is located in between $\epsilon_{s}$ and $\epsilon_{p_{-1}}$. This result is consistent with numerical results in the previous section; the coexistent nematic and singlet orders that break the translational symmetry sets in immediately when $\theta>0$, which corresponds to the S-phase.
@@
At $J_{2}<J_{1}/2$ in Figs. 8(a) and 8(b), the bottom of the band is located at $k=0$, while at $J_{2}>J_{1}/2$ (panels (c) and (d)), it shifts to $k=\pi$. It means that when $\theta$ reaches the phase boundary, the $p$-particles condense and form a magnetic long-range order of a period of the dimer (two sites) and two dimers (four sites), respectively. Indeed, the phase diagram obtained by DMRG in Fig. 1(a) has AF-UD2 and AF-UUDD phases, respectively, which are the states corresponding to S+N of different periodicity.
The bond operator approach provides both quantitatively and qualitatively nice agreement with the results of DMRG in the previous section. Indeed, the way the ND phase sustains at finite $\theta$ and how it differs from the AF-UUDD and UD phases or the TLL phases are clarified.
@@
Figure 9: Ground state phase diagram in an applied field $h$ at (a)-(d) $\theta=0.05\pi,0.1\pi,0.3\pi,0.5\pi$. We find ferromagnetism(F), up down(UD), flucuating-ferro(Fluct), vector chiral(VC), Tomonaga-Luttinger liquid (TLL),spin-density wave (SDW) phases, 1/3-plateau(P), and nematic-singlet dimer(ND).
@@
Figure 10: Two-point correlation functions $\langle O_{i}O_{j}\rangle$ of nematic operators $O_{i}=S^{x}_{i}S^{y}_{i+1}+S^{y}_{i}S^{x}_{i+1}$, vector chiral operators $O_{i}=S^{x}_{i}S^{y}_{i+1}-S^{y}_{i}S^{x}_{i+1}$, and $\langle O_{i}O_{j}\rangle-\langle O_{i}\rangle\langle O_{j}\rangle$ of spins $O_{i}=S_{i}^{\alpha_{i}}$ of $(\alpha_{i}\alpha_{j})=xx,zz$, and $xy$, obtained by DMRG with $N=100$. Symbols ND, VC, TLL, SDW, and Fluct-F indicate the nematic dimer, vector chiral, Tomonaga-Luttinger liquid, spin-density wave and fluctuating-Ferromagnetic phases, respectively. Several parameters are chosen from nonmagnetically ordered or fluctuating F phase.

### v2

@@ II.2.3 Correlation functions
We obtain the two-point correlation functions $\langle O_{i}O_{j}\rangle$, using the single-site or two-site operator $O_{i}$. For $O_{i}$, we take spin operators, $S_{i}^{\alpha}$ with $\alpha=x,y,z$, vector chiral operator $\kappa_{i}^{(n)}$, and nematic operator $Q^{xy}_{i}$. Figure 4 shows these correlation functions as functions of distance, $|i-j|$, obtained using SSD-DMRG. Let us summarize the features of each phase.
@@
TLL phase. Basically, the correlation functions all decay algebraically up to $|i-j|\leq 50$, where we reach half of the system size. However, at $\theta\neq 0$, $\langle Q_{i}^{xy}Q_{j}^{xy}\rangle$ shows a robust constant value at large distances, unlike the standard TLL phase. This is because if the system remains paramagnetic without any other orderings, the $\Gamma$ term works as a conjugate field to induce a finite value of $\langle Q_{i}^{xy}\rangle$, which is obvious from Eq.(1). In that context, the emergent $Q_{i}^{xy}$ is trivial as it does not break any symmetry of the Hamiltonian.
@@
Nematic-singlet dimer (ND) phase. Along the exactly solvable MG line $J_{2}/J_{1}=\Gamma_{2}/\Gamma_{1}=1/2$ at $\theta<{\rm arctan}(\sqrt{3})$, we find the dimer-product singlet states as an exact ground state, which strictly excludes other components and gives $\langle Q_{i}^{xy}Q_{j}^{xy}\rangle\rightarrow 0$. While, even when we are away from the MG line, the ND phase sustains where the magnetic correlation functions all decay exponentially, indicating the existence of a spin gap. Such gap opening is due to the breaking of translational symmetry in the same manner as the MG solution. The nematic correlation $\langle Q_{i}^{xy}Q_{j}^{xy}\rangle$ starts to converge to a small but finite constant value at long enough distances, indicating the formation of a long-range order with $\Gamma_{\eta}\neq 0$, as shown in Fig. 4, whose implication will be discussed shortly.
@@
Vector chiral (VC) phase. In the FM-AFM Heisenberg interaction range of the phase diagram at $\phi/\pi\sim 0.7-0.8$ ($J_{2}/J_{1}=-1.3\sim-0.72$), we find a VC phase where the magnetic order is absent and $\langle\kappa^{(1)}_{i}\kappa^{(1)}_{j}\rangle$ sustains at long distances. Its amplitude shows a large oscillation in the period of chirality over about $30$ spins as can be seen from the behavior of $\langle S_{i}^{x}S_{j}^{x}\rangle$ and $\langle S_{i}^{x}S_{j}^{y}\rangle$, while the pure magnetic components decay as a power law.
@@
Magnetically ordered phases. The FM phases, AFM-UD, and UUDD phases all have their magnetic easy axis pointing in the $x^{\prime}$ or $y^{\prime}$ directions perpendicular to the leg. Indeed, the correlation functions show a clear exponential decay about $\langle S^{z}_{i}S^{z}_{j}\rangle$, while those of the in-plane elements robustly take the constant value of order-1 throughout $|i-j|$. The nematic correlation is also robust because these magnetic orderings are the condensation of the off-diagonal $S=1$ elements.
II.2.4 Spin gap
@@
Our model does not conserve total-$S^{z}$, in which case the spin gap cannot be evaluated by the standard treatment of measuring the lowest-energy difference between different total-$S^{z}$ sectors. Instead, we apply a grand canonical appoach using SSDHotta and Shibata (2012); Hotta et al. (2013) that allows the access to the bulk magnetization curve for $N\gtrsim 20$ calculations (while we adopt $N=100$). By adding a Zeeman field $-h\sum_{j}S^{z}_{j}$ to Eq.(1), and by deforming them with the sine-square function, the magnetization $M$ is obtained by extracting the intrinsic values near the center using the $q=0$ element of the SSD Fourier transformationKawano and Hotta (2022). Figure 5(a) shows four different magnetization curves obtained for the nonmagnetic ground states, TLL, ND, VC, and N. Only the ND phase shows a substantially large spin gap, $\Delta\sim 0.3$. The spin gap of $J_{1}-J_{2}$ chain ($\Gamma=0$) was previously calculated by DMRGWhite and Affleck (1996) and was evaluated as $\Delta=0.11\sim 0.37$ for $0.14\pi<\phi<0.25\pi$ ($1.00.47<J_{2}/J_{1}<$ 1.0). For TLL and VC, the standard magnons condense and form a standard magnetization curve. Whereas, the magnetization curve of the N-phase exhibits a steep power-law increase starting from zero field, which reminds us of the magnetization curve of the ferrimagnetic-like stateYamashita et al. (2021). Such a high sensitivity of magnetization shall appear because the $S=1$ and $S^{z}=1$ gapless excitation from the N phase may exist.
@@
To confirm the presence of a finite spin gap, we examined the field-dependent magnetization in the direction of $-h\sum_{j}S^{\alpha}_{j}$, where $\alpha=x,y$, and $x^{\prime},y^{\prime}$ ($\pi/4$-rotation of $x,y$ about the $z$-axis) as shown in Appendix C; in the ND phase, the spin gap is finite and does not depend much on the direction of a field.
II.2.5 Nematic order parameter
@@
We have shown in Fig. 4 that the nematic correlation, $\langle Q^{xy}_{i}Q^{xy}_{j}\rangle$, starts to converge to a robust constant value once we introduce finite $\Gamma/J\neq 0$. To examine how they behave at around $\Gamma/J\sim 0$, we plot in Fig. 5(b) the averages of $\overline{\langle Q^{xy}_{i}Q^{xy}_{j}\rangle}$ over $|i-j|=58-61$ to exclude the strong oscillation effectHotta and Shibata (2012). Numerically, there is no “gap” in the onset value of $\big{(}\overline{\langle Q^{xy}_{i}Q^{xy}_{j}\rangle}\big{)}^{1/2}$, i.e. it increases immediately from $\Gamma=0$ in power of $\Gamma$ for all displayed parameters of $\phi$. As mentioned, $\Gamma$ in Eq.(1) works as chemical potential of nonmagnetic quasi-particle represented by the operator $Q^{xy}_{i}$.
@@
The AFM-UUDD has the staggard magnetic moment pointing in the $y^{\prime}$ direction which is nothing but the “magnetization” represented by $Q^{xy}$. Namely, the doped quasi-particles condense and form a regular four-fold periodic structure breaking the translational symmetry. In the ND phase, the two-fold periodic breaking of translational symmetry occurs but there is no magnetic order because of the interplay of singlet and nematic particles. The spin gap is open, which is distinct from the AFM-UUDD and TLL phases.
@@
For these reasons, we understand that “field” effect, $\Gamma$, generates a finite nematic correlation but it appears in different ways depending on the degree of frustration dictated by $\phi$. As we see shortly in §.II.3, the ND phase hosts gapless excitation mode due to $\Gamma$ that gives the finite but small distribution of $\langle Q^{xy}_{i}\rangle$ in real space on top of the singlets, and at large $\Gamma$ the nematic particles finally condense into the magnetically ordered phases, AFM-UD2, AFM-UUDD. The TLL phase, although hosting finite nematic correlation, shall be trivial.
Figure 6: (a,b) Spin and (c-f) nematic dynamical structure factors, ${\cal S}_{m}(q,\omega)$ and ${\cal S}_{Q}(q,\omega)$. For the density plots given in (a-d), we take $\phi/\pi=0.2$ corresponding to $J_{2}/J_{1}\approx 0.73$ with $\theta/\pi=0$ ($\Gamma=0$) and $0.1$($\Gamma/J\approx 0.32$) for the pure singlet dimer and ND phases, respectively. (e) The frequency dependences of ${\cal S}_{Q}$ are shown for fixed wave numbers, $q=\pi/2$ (solid line) and $\pi$ (broken line), where we take $(\phi/\pi,\theta/\pi)=(0.2,0),(0.2,0.1),(0.12,0.1)$. The green and blue arrows indicate the $q=\pi/2$ and $q=\pi$ peaks at $\omega\sim 0$, respectively. (f) The density plot of ${\cal S}_{Q}$ at $(\phi/\pi,\theta/\pi)=(0.12,0.1)$ ($J_{2}/J_{1}\approx 0.40,\Gamma/J\approx 0.32$) is shown which is to be compared with panel (d).
II.3 Dynamical structure factor
As we found in Fig. 5(b), the long-distant value of $\langle Q^{xy}_{i}Q^{xy}_{j}\rangle$ increases in power of $\Gamma/J$, suggesting that the nematic order parameter $\langle Q^{xy}\rangle\sim\sqrt{\langle Q^{xy}_{i}Q^{xy}_{j}\rangle}$ becomes finite by the introduction of infinitesimally small $\Gamma/J$, even for the ND phase where the singlet long-range order is present. When regarding $\Gamma$ as a “field” coupled to $Q^{xy}$, this indicates that the ground state is gapless. To confirm it, we calculate the dynamical structure factors using the standard time-evolving block decimation (TEBD) techniqueVidal (2004) as,
${\cal S}_{m/Q}(q,\omega)$ $=\frac{1}{N}\sum_{j=1}^{N}\int\frac{dt}{2\pi}e^{i(\omega t-qr_{j})-\eta^{2}t^{2}}G_{m/Q}(r_{j},t),$ (8)
$G_{m}(r_{j},t)=\;\langle 0|\bm{S}_{j}(t)\cdot\bm{S}_{N/2}(0)|0\rangle\;,$
$G_{Q}(r_{j},t)=\;\langle 0|Q^{xy}_{j}(t)\cdot Q^{xy}_{N/2}(0)|0\rangle,$ (9)
where $|0\rangle$ the ground state of the Hamiltonian and $r_{j}=j-N/2$ is the one-dimensional coordinate of site $j$ measured from the center site $N/2$. We consider two types of dynamical structure factors, ${\cal S}_{m}(q,\omega)$ and ${\cal S}_{Q}(q,\omega)$, that takes account of the magnon and nematic excitationsRamos et al. (2018); Tóth et al. (2012), respectively. The nematic excitations are indeed detected in resonant inelastic x-ray scattering (RIXS)Haverkort (2010); Ament et al. (2011). Notice that since we take the lattice spacings between $j$ and $j+1$ as a unit in the Fourier transformation the reciprocal number $q$ is defined accordingly.
We prepare an initial state $\bm{S}_{N/2}|0\rangle$ and $Q^{xy}_{N/2}|0\rangle$ and perform the TEBD with a timestep of $\delta t=0.1$ up to the maximum time, $T_{{\rm max}}=34\sim 37$ for $N=400$ with open boundary using the maximum bond dimension, $\chi=400$, and the Gaussian broadening $\eta^{2}=0.004$.
Figures 6(a)-(d) show ${\cal S}_{m}$ and ${\cal S}_{Q}$ in the pure dimer $(\phi/\pi,\theta/\pi)=(0.2,0)$ and nematic-singlet dimer $(0.2,0.1)$ phases. The spin dynamical structure factors in panels (a,b) clearly show a finite gap $\Delta/J_{1}\sim 0.2-0.3$ and the peak wave number $q\sim 0.7,1.3$, which is consistent with the spin gap in Figure 5(a) and the incommensurate period $q_{\text{peak}}$ in Figure 3(d), demonstrating that the single magnon dispersion is insensitive to $\Gamma$ terms. A strong peak structure is observed near $\omega=\Delta$ at $q=q_{\text{peak}}$, which is consistent with the results for the Heisenberg zigzag ladder in Ref.[Sacramento and Vieira, 2001] at $J_{2}/J_{1}=1,\Gamma/J=0$.
Regarding the nematic dynamical structure factor in Figs. 6(c-f), there is a the distinct difference between those of zero and finite $\Gamma$; When $\Gamma=0$ in panel (c), the spectrum has a gap up to $\omega\sim J_{1}$, whereas introducing $\Gamma\neq 0.1$ in panel (d), there appears a new weight at $\omega\sim 0$ and $q=\pi/2$ indicated by an arrow. In Fig. 6(e), the intensity of ${\cal S}_{Q}(\omega,q)$ for fixed $q$ is given as a function of $\omega$, where the $\omega=0$ peak is found to be robust. We confirmed that the peak positions are located at $q=\pi$ for $J_{2}/J_{1}<0.5$ and at $q=\pi/2$ for $J_{2}/J_{1}>0.5$, which are consistent with the ordering periods of AFM-UD2 and UUDD phases, respectively, that appear in the larger $\Gamma$-part of the phase diagram.
We thus find that there is a substantial gapless low energy component carried by $\langle Q^{xy}\rangle$, which is less dispersive and conforms to the translational symmetry broken structure of single dimers. This gapless weight is the precursor for the magnetic long-range orderings that take place at large $\Gamma/J$.
@@ III Bond-operator mean-field theory
In this section, we examine the effect of $\Gamma$-term in the AFM-AFM zigzag chain ($J_{1},J_{2}>0,0\leq\phi\leq\pi/2$) by using bond-operator mean-field theorySachdev and Bhatt (1990). We see that in the Hamiltonian Eq.(1) $\Gamma$ works as a “field” (or a chemical potential) to condense the nematic particle represented by $Q_{i}^{xy}$. This situation is similar to doping a magnon to the gapped singlet stateSachdev and Bhatt (1990). The condensation of nematic particles is discussed previously, where a ring exchange interaction or frustrated interactions work as a chemical potentialUeda and Totsuka (2007); Totsuka et al. (2012); Yokoyama and Hotta (2018). All these cases take a gapped singlet ground state as a starting point, which is approximated by the product state of dimers (or tetramers), and the effect of the applied “field” is examined using the bond-operator approach or its analog.
@@ III.2 Effect of inter-dimer interaction
$S_{2}^{\alpha}$ $=\frac{1}{2}(-s_{n}^{\dagger}p_{\alpha,n}-p_{\alpha,n}^{\dagger}s_{n}-i\epsilon_{\alpha\beta\gamma}p_{\beta,n}^{\dagger}p_{\gamma,n}),$ (13)
@@
Here, we consider a small $\lambda$, assuming $\gamma<2$ ($\theta<0.35\pi$), expecting that the ground state is dominated by singlets and the population of $p_{\alpha,n}$-bosons are dilute. The condensed singlets are replaced by its expectation value $\langle s\rangle=\bar{s}$, and by dropping off the terms consisting of three or four $p$ operators, we obtain the interactions between the $n$-th and $(n+1)$-th dimers as
@@ III.4 Phase diagram
The phase diagrams obtained by the bond-operator mean-field theory are shown in Figs. 7(b) and 7(c) for several fixed values of $\lambda=0.1,0.5,1.0$ and $\phi/\pi=0.08$ to $0.4$ ($J_{2}/J_{1}=0.26-3.1$), respectively. The corresponding set of mean-field parameters $(\mu,\bar{s}^{2})$ are shown in Fig. 8 at $\lambda=1$ as functions of $\theta$. For these parameters, there is a range of small $\theta$ that the ground state is fully occupied by singlets, $\bar{s}^{2}=1$, which we call S-phase. The phase transition from S-phase to (S+N)-phase takes place at finite $\theta$, where $\bar{s}^{2}$ starts to decrease from 1, and the chemical potential $\mu$ reaches the singlet (zero) energy level. When increasing $\lambda$, the S-phase shrinks because the $p$-particles gain the kinetic energy due to inter-dimer interactions, but at $\phi={\rm atan}(1/2)$, i.e. $J_{2}/J_{1}=\Gamma_{2}/\Gamma_{1}=0.5$, which corresponds to the Majumdar-Ghosh line, the singlet phase remains stable throughout the whole range of $\theta$. This fact is trivial as we find that the inter-dimer interaction in Eq.(14) is $\hat{h}_{n}^{\rm MF}=0$, or equivalently, the dispersion in Eq.(18) yields $\epsilon(k)=0$ for $J_{1}=2J_{2}$. Indeed, along this line, the singlet product state is the exact solution of the original Hamiltonian Eq.(1) without any approximation.
@@
We plot in Figs. 9(a)-9(d) the energy dispersions $\omega_{l}(k)$ at $\lambda=1$ for several choices of $\phi$ and $\theta$, inside the S-phase. Remarkably, all of them are gapless. Indeed, the energy gap $\Delta$ of ${\cal H}_{\rm bo}$ shown in Fig. 9(e) is zero for all parameters of $\phi$ at small $\theta$. This shows that the singlet phase is robust but at the same time, the $p$-particle are always able to join the ground state. This result is rather unusual; the standard bond operator approach is used to provide the instability toward the condensation of these excited $p$-state that replaces the ground state, and the gap-closing point usually indicates the phase transition. Whereas, in the present case, the zero gap of ${\cal H}_{\rm bo}$ does not mean the instability of the singlet ground state. We emphasize that the instability takes place not at the gap closing point but when $\bar{s}^{2}$ starts to deviate from 1, and the chemical potential $\mu$ reaches the singlet energy level. For clarification, we plot in Fig. 8(b) together the bare excited energy, $\epsilon_{p_{-1}}$, of the isolated dimer. We see that $\mu$ is located in between $\epsilon_{s}$ and $\epsilon_{p_{-1}}$. This result is consistent with numerical results in the previous section; the coexistent nematic and singlet orders that break the translational symmetry sets in immediately when $\theta>0$, which corresponds to the S-phase.
@@
The $\omega_{2}$ branch in Fig. 9 has a gapless point at $k=0$ ((a,b) $J_{2}/J_{1}<0.5$) and $k=\pi$ ((c,d)$J_{2}/J_{1}>0.5$) for periods of dimer. They agree with the gapless nonzero weight of the dynamical structure factor $q=\pi$ ($J_{2}/J_{1}<0.5$) and $q=\pi/2$ ($J_{2}/J_{1}>0.5$) defined in the unit of lattice sites in Fig. 6, respectively. When $\theta$ reaches the phase boundary, the whole $\omega_{2}$ branch condenses and replaces the singlet, and the magnetic long-range order in a period of the dimer (two sites) and two dimers (four sites) appear for $J_{2}/J_{1}<0.5$ and $>0.5$, respectively. Indeed, the phase diagram obtained by DMRG in Fig. 1(a) has AFM-UD2 and AFM-UUDD phases, respectively, which are the states corresponding to S+N of different periodicity.
We may interpret that the dispersiveness of $\omega_{2}$ around the gapless point shall be the artifact of the mean-field approximation to one-body because the higher order interactions and many-body effects are discarded (see Appendix D). Due to the correlation effects the weight of dispersions off the gapless point shall be weakened to form the gapless peak in the dynamical structure factor in Fig. 6.
The $p_{1,n}^{\dagger}$ and $p_{-1,n}^{\dagger}$ are the creation operator on the $n$-th dimer in Eq.(12) of the states that give $\langle Q^{xy}_{n}\rangle=1$ and $-1$, respectively. Therefore, we evaluated the number operators with respect to the eigenstate $|k,2\rangle$ of $\omega_{2}(k)$, finding that
$\langle k,2|p_{1,n}^{\dagger}p_{1,n}|k,2\rangle=0,$ $\langle k,2|p_{-1,n}^{\dagger}p_{-1,n}|k,2\rangle=2/N,$ (19)
for all $k$. This indicates that each $k$-point carries a single $p_{-1,n}$ particle. The gapless excitation that contributes to the ground state as fluctuations thus yield the dilute order-$1/N$ concentration of $p_{-1,n}$, in agreement with the DMRG value of $\langle Q^{xy}_{i}\rangle\sim 0.01$ in Figs. 4(d) 5(b).
@@
Figure 10: Ground state phase diagram in an applied field $h$ at (a)-(d) $\theta=0.05\pi,0.1\pi,0.3\pi,0.5\pi$ ($\Gamma/J\approx 0.16,0.32,1.4,\infty$). We find ferromagnetism(FM), up down(UD), fluctuating-ferro(Fluct), vector chiral(VC), Tomonaga-Luttinger liquid (TLL),spin-density wave (SDW) phases, 1/3-plateau(P), and nematic-singlet dimer(ND).
@@
Figure 11: Two-point correlation functions $\langle O_{i}O_{j}\rangle$ of nematic operators $O_{i}=S^{x}_{i}S^{y}_{i+1}+S^{y}_{i}S^{x}_{i+1}$, vector chiral operators $O_{i}=S^{x}_{i}S^{y}_{i+1}-S^{y}_{i}S^{x}_{i+1}$, and $\langle O_{i}O_{j}\rangle-\langle O_{i}\rangle\langle O_{j}\rangle$ of spins $O_{i}=S_{i}^{\alpha_{i}}$ of $(\alpha_{i}\alpha_{j})=xx,zz$, and $xy$, obtained by DMRG with $N=100$ with $\theta=0.05\pi$ ($\Gamma/J\approx 0.16$). Symbols ND, VC, TLL, SDW, and Fluct-FM indicate the nematic dimer, vector chiral, Tomonaga-Luttinger liquid, spin-density wave and fluctuating-Ferromagnetic phases, respectively. Several parameters are chosen from nonmagnetically ordered or fluctuating FM phase.